# Remote Sensing Reveals Adoption of Sustainable Rice Farming Practices Across Punjab, India

### v1

@@ Abstract
Rice cultivation consumes 24-30% of global freshwater, creating critical water management challenges in major rice-producing regions. Sustainable irrigation practices like direct seeded rice (DSR) and alternate wetting and drying (AWD) can reduce water use by 20-40% while maintaining yields, helping secure long-term agricultural productivity as water scarcity intensifies—a key component of the Zero Hunger Sustainable Development Goal. However, limited data on adoption rates of these practices prevents evidence-based policymaking and targeted resource allocation. We developed a novel remote sensing framework to monitor sustainable water management practices at scale in Punjab, India—a region facing severe groundwater depletion of 41.6 cm/year. To collect essential ground truth data, we partnered with the Nature Conservancy’s Promoting Regenerative and No-burn Agriculture (PRANA), which trained approximately 1,400 farmers on water-saving techniques while documenting their field-level practices. Using this data, we created a classification system with Sentinel-1 satellite imagery that separates water management along sowing and irrigation dimensions. Our approach achieved a 78% F1-score in distinguishing DSR from traditional puddled transplanted rice without requiring prior knowledge of planting dates. We demonstrated scalability by mapping DSR adoption across approximately 3 million agricultural plots in Punjab, with district-level predictions showing strong correlation (Pearson=0.77, RBO= 0.77) with government records. This study provides policymakers with a powerful tool to track sustainable water management adoption, target interventions, and measure program impacts at scale.
Index Terms—  Deep Learning, Remote Sensing, GeoAI, Rice Agriculture, Applied Machine Learning, Sustainable Agriculture, Water Management
@@ 1 Introduction
Rice is a staple food for half of the world’s population, but traditional paddy cultivation requires enormous quantities of water—2,000–5,000 liters per kilogram of rice produced [1]. As climate change intensifies water scarcity in many regions, sustainable water management in rice cultivation becomes essential for ensuring long-term food security, critical to addressing the U.N. Zero Hunger Sustainable Development Goal (SDG). Two water-saving practices have emerged as particularly promising: Direct Seeded Rice (DSR) and Alternate Wetting and Drying (AWD). DSR eliminates the need to raise and transplant seedlings by directly sowing seeds into the field, reducing water consumption by 20-40% [2, 1], relative to traditionally practiced puddled transplanted rice (PTR). AWD involves periodically allowing fields to dry before re-flooding, reducing water use by up to 30% compared to traditional continuous flooding (CF) [3]. When properly managed, these techniques maintain or even enhance yields, presenting a balanced approach for sustainable rice production [4].
@@
Fig. 1: Mapping DSR predictions across Punjab for 2024: A) Map of Punjab showing DSR adoption distributions with district boundaries. B) Comparison between model predictions and government records of DSR adoption by district, ordered by acres of DSR activity, as estimated by the Punjab government. C) Correlation analysis (Pearson=0.79) between model predictions and government records with outlier and important districts labeled.
@@
Remote sensing offers a potential solution for monitoring water management practices at scale. While satellite-based methods have successfully mapped rice cultivation using multispectral optical and synthetic aperture radar (SAR) approaches [5, 6, 7], detecting specific water management practices remains challenging. Existing studies rely heavily on prior knowledge of planting dates [8, 9] or use coarse resolution data [10], limiting their scalability in regions where planting dates vary significantly and in regions dominated by smallholder farming.
@@
We focused our analysis on Punjab, India, a major rice-producing region facing severe water stress. Punjab contributes 10.42% of India’s rice production [11], but faces groundwater depletion of 41.6 cm/year [12]. This crisis stems from the region’s semi-arid climate, rice-wheat cropping systems that increased irrigation demand, and groundwater extraction facilitated by free electricity [13, 14]. Approximately 73% of irrigation water comes from rapidly depleting groundwater sources [15]. To address these challenges, we partnered with The Nature Conservancy’s Promoting Regenerative and No-burn Agriculture (PRANA) project [^1] https://www.nature.org/en-us/about-us/where-we-work/india/our-priorities/prana/. PRANA trained approximately 150,000 farmers across Punjab on sustainable water management methods, while collecting detailed field-level data from about 1,400 participants. This included documentation of sowing methods, irrigation practices, and precise field boundaries–critical ground truth data for our remote sensing approach. Our study introduces a dimensional framework for monitoring water management practices using Sentinel-1 satellite imagery. By separating practices along sowing (DSR vs. PTR) and irrigation (AWD vs. CF) dimensions, we leverage their distinct temporal signatures while avoiding the limitations of simultaneous detection. Importantly, our approach does not require prior knowledge of cropping calendars, enhancing its scalability for large-scale monitoring.
Fig. 2: Water-saving Interventions: A) AWD measuring pipe being installed in a demonstration field. B) AWD pipe after installation. C) Photograph of a sample field practicing DSR taken during the tillering stage; note the lack of flooding in the field.
Our main contributions include:
  • A dimensional classification framework that separates water management practices along sowing and irrigation dimensions
  • A robust model using Sentinel-1 time-series imagery that achieves 78% F1-score in classifying DSR despite significant variability in planting dates
@@
  • A scalable plot-level method demonstrated across approximately 3 million agricultural plots in Punjab, with aggregated district-level predictions correlating strongly with government records
@@ 2 Background
2.1 Study Area
Our study focuses on Punjab, India (Fig. 5), a semi-arid region where rice dominates agriculture, covering 87% (3.13 million hectares) of farmland during the 2022-23 monsoon season [16]. In 2023, Punjab produced 14.36 million tons of rice—10.4% of India’s production and approximately 3.3% of global output [11]. The region typically employs a wheat-rice double-cropping system using PTR and CF methods[17].
This agricultural intensity has created a severe water crisis, with 73% of irrigation water drawn from rapidly depleting underground aquifers. The 2022 Central Ground Water Board report classified 114 of Punjab’s 150 blocks as over-exploited, with an additional 19 blocks categorized as critical or semi-critical [18]. Although the state government has implemented conservation measures, including legislation restricting early rice transplanting [15], groundwater levels continue to decline at an alarming rate of 41.6 cm/year, expected to drop below 300 metres by 2039 [12, 18].
Fig. 3: Temporal view of water management practices. Y-axis represents water level in a rice paddy, and x-axis shows growth stages. The growing season naturally separates into two phases: initial (left of dotted line) and main growth (right). Blue arrows indicate irrigation events, while yellow arrows denote sowing. The different practices create distinct temporal signatures, with the vertical dotted line marking the boundary between sowing and irrigation dimensions.
Fig. 4: Dimensional divisions of practices. A) Plots divided by sowing and irrigation dimensions, with number of plots in parentheses. B) Classification of irrigation practices (grouping Control and DSR as CF). C) Classification of sowing practices (grouping Control and AWD as PTR). Dotted red lines in B) and C) show classes are separated in a task-dependent manner.
2.2 Challenges
@@
2.2.1 Water Management Practices and Separability
Water-saving practices in rice cultivation occur at different crop stages and create distinct water level patterns. DSR affects water management during early sowing and tillering phases, while AWD impacts water dynamics from tillering through harvest.
As shown in Figure 4, these practices create unique temporal signatures in water levels and soil moisture that can be detected via Synthetic Aperture Radar (SAR) remote sensing [19, 8, 9, 10]. PTR maintains consistently high water levels, while DSR shows a sharp initial decline after pre-sowing irrigation. Similarly, CF maintains flooded fields throughout the season, whereas AWD creates cyclical wet-dry patterns.
2.2.2 Dimensional Classification Approach
@@
This dimensional approach, illustrated in Figure 4, allows us to cleanly separate practices and collect training data that isolates specific dimensions of interest. Control plots employ PTR and CF practices, DSR plots practice DSR and CF, and AWD plots practice PTR and AWD. Plots practicing mixed combinations were excluded to avoid confounding effects.
2.2.3 Planting Date Heterogeneity
An additional challenge for satellite-based monitoring is the significant variation in planting schedules. Previous studies have relied on synchronizing satellite data with known planting dates [8, 9] – a luxury unavailable when monitoring millions of plots. In our study area, planting dates span up to 110 days, with the majority of DSR plots typically sown first, followed by control plots and finally AWD plots (Fig. S3). This temporal variability introduces natural differences between classes that must be accounted for, and in fact is exploited, in our modeling approach.
2.2.4 Plot-level Predictions
To be able to provide fine-grained insights, we must be able to both identify individual plot boundaries as well as make predictions about these monolithic entities. We employ Fields of the World (FTW) algorithm [20] to automate the extraction of plot boundaries, enabling us to make predictions at an individual plot level.
Our research question thus becomes: How effectively can we classify water management practices by separating them along sowing and irrigation dimensions using satellite imagery, despite significant heterogeneity in planting dates, while predicting at the plot-level?
@@ 3 Materials and Methods
3.1 Ground Truth Data Collection
To obtain reliable training data, we partnered with The Nature Conservancy’s PRANA project, which trains farmers on sustainable water management practices across Punjab. Between May and December 2024, local field teams collected detailed plot-level data from PRANA intervention sites across 18 districts. The resulting dataset includes 452 AWD plots, 420 DSR plots and 411 PTR plots employing CF. Their average sizes are 4.2, 2.7 and 2.1 acres, respectively. Refer to Fig. S1 for details of AWD measuring devices and visuals of DSR plotsin the PRANA program.
For each plot, we documented sowing dates, transplantation dates (for PTR/AWD), irrigation schedules, crop varieties, harvesting dates, and precise field boundaries. Field delineations were conducted through on-site visits using custom software and Google Earth Pro. We selected only plots with clearly visible boundaries and a minimum area of 2000 square meters to ensure sufficient quality in satellite data analysis. A 10-meter negative buffer was applied to all boundaries to eliminate edge effects while preserving adequate pixels for analysis.
3.2 Satellite Data and Processing
We leveraged Sentinel-1 Synthetic Aperture Radar (SAR) data to monitor water management practices across the growing season (April-December 2024). We acquired 108 images from ascending orbit — 17 $\times$ 3 from row 27222An on-board anomaly on Sentinel-1A on June 27-28 led to irrecoverable loss of data and therefore a missing timestep in the row 27 and 19 $\times$ 3 from row 100 — covering the entire state of Punjab. Through experimentation, the ascending orbit dataset was shown to have higher performance and was selected for training and inference, focusing on the time period May 1 - December 15, the maximum duration of the kharif growing season for rice in Punjab. For details please refer to section 4.2.
Fig. 5: Map of Punjab showing the study area, Sentinel-1 satellite coverage (ascending and descending orbits), and locations of surveyed plots. Insets show sample plots at different zoom levels, with the innermost inset displaying delineated fields from three classes: Control (plots that use puddled transplanting and continuous flooding), alternate wetting and drying (AWD), and direct seeded rice (DSR).
SAR data was specifically selected for its cloud-penetrating capabilities, a crucial advantage during the monsoon season when optical imagery is frequently compromised. Previous have shown SAR’s effectiveness in capturing diverse rice agriculture metrics, including cultural type classification, phenological parameter detection, growing season estimation [21, 22, 23], and differentiation between multiple cropping cycles and field-level tillage practices [7, 24]. To account for environmental factors influencing irrigation decisions, we integrated ERA5-Land climate reanalysis data [25], incorporating daily temperature and precipitation values as contextual variables in our analysis framework.
All SAR imagery was processed using the SNAP toolbox [26] to extract time series information at 10-meter spatial resolution. We extract Sentinel-1 VV and VH polarization bands from multiple orbit passes over Punjab. For each image, we apply standard preprocessing steps: orbit file application, noise removal, radiometric calibration, speckle filtering using multi-temporal refined-Lee filter [27], and terrain correction to derive terrain-geocoded backscatter coefficient ($\gamma_{0}$) at 10-meter resolution [26].
@@ 3.2 Satellite Data and Processing
For each agricultural plot polygon in the training corpus, we calculate mean backscatter values for every timestep, generating a timeseries per plot. This process serves both as speckle noise reduction for the SAR imagery, and as a summary of plot-level conditions. To account for temporal gaps between satellite passes and enable consistent analysis, we apply temporal smoothing with a 1-D Gaussian filter (standard deviation = 0.5) followed by fitting a third-degree polynomial spline. This allows us to sample the time series at regular intervals despite variable satellite acquisition dates–steps that are crucial in dealing with missing data in the timeseries. We additionally create a VV/VH band ratio time series, which has proven effective for rice mapping in previous studies [9, 28, 29].
3.3 Feature Engineering
We develop two complementary sets of features:
@@ 3.3 Feature Engineering
Hand-crafted Features: For each band (VV, VH, VV/VH), we extract temporal characteristics that include the position and amplitude of the first $k$ troughs, crests, and inflection points; the total number of troughs and crests; summary statistics (mean, minimum, maximum amplitude); parameters from Gaussian kernel fitting to the VV/VH ratio (amplitude, peak day, standard deviation and $R^{2}$)[28]; and radar vegetation index (RVI) values [30].
Pretrained Features: We leverage the Pretrained Remote Sensing Transformer (Presto) [31], a deep learning model pre-trained on diverse remote sensing time series data. We picked this remote sensing foundation model as it works at a pixel level as opposed to patch level, which would be difficult due to the irregular size of plots. We extract embeddings by passing ERA5 climate variables and Sentinel-1 timeseries data through the frozen Presto encoder, generating features that capture complex temporal patterns without requiring extensive labeled data. We utilize the embeddings from all timesteps (“temporal tokens”). A limitation of this model is that it can process at most 24 timesteps. We experiment with resampling the smoothed spline in the previous step with frequencies of 4, 7 and 10 days. This is done in order to increase the temporal resolution of the data where possible (when the total number of timesteps $\leq$ 24) or to accommodate longer temporal ranges (such as rows 11-12 in Table 2).
3.4 Classification Approach
@@
We test two ensemble learning methods—Random Forest (RF) and Gradient Boosted Trees (GB)—for their proven effectiveness in agricultural classification tasks [32, 33]. These models effectively handle the heterogeneous feature sets while remaining computationally efficient. Following our dimensional framework, we structure three classification tasks: a) AWD vs. DSR vs. Control (null hypothesis: all practices on same dimension), b) DSR vs. PTR (sowing dimension), and c) AWD vs. CF (irrigation dimension). For each task, we split data into non-overlapping train/test sets (90:10 ratio) using stratified sampling. We optimize hyperparameters using Optuna [34] and evaluate performance using overall accuracy and F1 scores (macro and globally weighted averages). We compare against a baseline that randomly assigns classes based on class distribution proportions. This approach tests our hypothesis that separating water management practices along their natural dimensions (sowing and irrigation) produces more accurate results than attempting simultaneous multi-class classification (Fig. S1 for the complete workflow).
Classification Task Model Overall Accuracy F1 (weighted) F1 (macro)
AWD v/s DSR v/s PTR Baseline 33.47% (±3%) 33.38% (±3%) 33.27% (±3%)
RF 55.04% 54.39% 54.11%
GB 59.69% 59.62% 59.57%
DSR v/s PTR Baseline 56.8% (±3%) 56.3% (±3%) 50.01% (±3%)
RF* 75.19% 75.19% 71.76%
GB* 77.52% 77.82% 75.13%
AWD v/s CF Baseline 53.92% (±3%) 53.89% (±3%) 49.98% (±3%)
RF** 79.84% 78.76% 75.75%
GB** 80.62% 80.33% 78.07%
Table 1: Classification results for dimensional and combined approaches for the best performing models. RF = Random Forest, GB = Gradient Boosted Trees. Bold indicates best results per task. *=sampling frequency of 4 days, **=sampling frequency of 10 days

### v2

# Machine Learning for Sustainable Rice Production: Region-Scale Monitoring of Water-Saving Practices in Punjab, India

@@ Abstract
Rice cultivation supplies half the world’s population with staple food, while also being a major driver of freshwater depletion–consuming roughly a quarter of global freshwater–and accounting for $\sim$48% of greenhouse gas emissions from croplands. In regions like Punjab, India, where groundwater levels are plummeting at 41.6 cm/year, adopting water-saving rice farming practices is critical. Direct-Seeded Rice (DSR) and Alternate Wetting and Drying (AWD) can cut irrigation water use by 20–40% without hurting yields, yet lack of spatial data on adoption impedes effective adaptation policy and climate action. We present a machine learning framework to bridge this data gap by monitoring sustainable rice farming at scale. In collaboration with agronomy experts and a large-scale farmer training program, we obtained ground-truth data from $\sim$1,400 fields across Punjab. Leveraging this partnership, we developed a novel dimensional classification approach that decouples sowing and irrigation practices, achieving F1 scores of 0.8 and 0.74 respectively, solely employing Sentinel-1 satellite imagery. Explainability analysis reveals that DSR classification is robust while AWD classification depends primarily on planting schedule differences, as Sentinel-1’s 12-day revisit frequency cannot capture the higher frequency irrigation cycles characteristic of AWD practices. Applying this model across 3 million fields reveals spatial heterogeneity in adoption at the state level, highlighting gaps and opportunities for policy targeting. Our district-level adoption rates correlate well with government estimates (Spearman’s $\rho$=0.69 and Rank Biased Overlap=0.77). This study provides policymakers and sustainability programs a powerful tool to track practice adoption, inform targeted interventions, and drive data-driven policies for water conservation and climate mitigation at regional scale.
Code & Datasethttps://github.com/microsoft/rice-irrigation-mapping-s1s2
Extended versionhttps://arxiv.org/abs/2507.08605
@@ 1 Introduction
Rice is a staple food for half of the world’s population, but traditional paddy cultivation requires enormous quantities of water—2,000–5,000 liters per kilogram of rice produced (bouman2007water), consuming 24-30% of world’s freshwater (surendran2021use). As climate change intensifies water scarcity in many regions, sustainable water management in rice cultivation becomes essential for ensuring long-term food security, critical to addressing the United Nations Zero Hunger Sustainable Development Goal (SDG). Furthermore, rice cultivation accounts for $\sim$48% of greenhouse gas (GHG) emissions from croplands as a direct consequence of water-intensive paddy cultivation (qian2023greenhouse). Two water-saving practices have emerged as particularly promising: Direct Seeded Rice (DSR) and Alternate Wetting and Drying (AWD). DSR eliminates the need to raise and transplant seedlings by directly sowing seeds into the field, reducing water consumption by 20-40% (bhushan2007saving; bouman2007water), relative to traditionally practiced puddled transplanted rice (PTR). AWD involves periodically allowing fields to dry before re-flooding, reducing water use by up to 30% compared to traditional continuous flooding (CF) (richards2014alternate). When properly managed, these techniques maintain or even enhance yields, presenting a balanced approach for sustainable rice production (lansing2023adaptive).
@@
[…]
@@
Remote sensing offers a potential solution for monitoring water management practices at scale. While satellite-based methods have successfully mapped rice cultivation using multispectral optical and synthetic aperture radar (SAR) approaches (nguyen2016mapping; fatchurrachmanHighResolutionMappingPaddy2022; singhaHighResolutionPaddy2019), detecting specific water management practices remains challenging. Existing studies rely heavily on prior knowledge of planting dates (fikriyahDiscriminatingTransplantedDirect2019; villano2019separability) or use coarse resolution data (gumma2015mapping), limiting their scalability in regions where planting dates vary significantly, or smallholder-dominated regions.
@@
We focused our analysis on Punjab, India, a major rice-producing region facing severe water stress. Punjab contributes 10.42% of India’s rice production (MARKETINTELLIGENCEREPORT2024), but faces groundwater depletion of 41.6 cm/year (baweja2017groundwater). This crisis stems from the region’s semi-arid climate, rice-wheat cropping systems that increased irrigation demand, and groundwater extraction facilitated by free electricity (gupta2023free; PunjabStateAgriPolicy2023). Approximately 73% of irrigation water comes from rapidly depleting groundwater sources (sidhu2021spatio). To address these challenges, we leveraged data from The Nature Conservancy’s Promoting Regenerative and No-burn Agriculture (PRANA) project [^1] https://www.nature.org/en-us/about-us/where-we-work/india/our-priorities/prana/. PRANA trained approximately 150,000 farmers across Punjab on sustainable water management methods, while collecting detailed field-level data from about 1,400 participants. This included documentation of sowing methods, irrigation practices, and precise field boundaries–critical ground truth data for our remote sensing approach. Our study introduces a dimensional framework for monitoring water management practices using Sentinel-1 (S1) satellite imagery. By separating practices along sowing (DSR vs. PTR) and irrigation (AWD vs. CF) dimensions, we leverage their distinct temporal signatures while avoiding the limitations of simultaneous detection. Conceptually, this framing clarifies that agronomic processes driving sowing and irrigation decisions are distinct and temporally separable. Importantly, our approach does not require prior knowledge of cropping calendars, enabling large-scale monitoring, with results shown in Figure 1. Our main contributions include:
  • Methodologically, we develop a comparative evaluation framework that operationalizes this decomposition at state scale, integrating features extracted from S1 using traditional methods and pretrained Earth Observation (EO) models.
  • Empirically, we show that this formulation achieves higher accuracy than a three-class model (0.798 v/s 0.616). We conduct explainability analysis of model outputs to provide evidence that high irrigation classification performance (0.742 F1) stems from planting schedule artifacts rather than irrigation patterns of frequencies higher than can be currently captured by a single S1 satellite.
@@
  • Practically we demonstrate scalable and near-operational monitoring of >>3 million agricultural plots in Punjab, with aggregated district-level predictions correlating strongly with government records.
@@ 2 Background
Study Area
Figure 2: Map of Punjab showing the study area, Sentinel-1 satellite coverage (ascending and descending orbits), and locations of surveyed plots. Insets show sample plots at different zoom levels, with the innermost inset displaying delineated fields from three classes: Control (plots that use puddled transplanting and continuous flooding), alternate wetting and drying (AWD), and direct seeded rice (DSR). Background satellite imagery from Bing Maps.
Our study focuses on Punjab, India (Fig. 2), a major rice-producing region facing severe water stress. Punjab contributes 10.42% of India’s rice production and approximately 3.3% of global output (MARKETINTELLIGENCEREPORT2024), but faces rapid groundwater depletion. This crisis stems from wheat-rice double-cropping system using PTR and CF methods (han2022annual) and groundwater extraction facilitated by free electricity, with 73% of irrigation extracted from rapidly depleting underground aquifers. A 2022 Government report classified 114 of Punjab’s 150 blocks as over-exploited (CGWBreport2022). To address these challenges, we partnered with The Nature Conservancy’s PRANA project, which trained approximately 150,000 farmers on sustainable water management while collecting detailed field-level data from more than 1,400 participants.
Figure 3: Water management practices across classes, created in collaboration with rice agronomy experts. The $y$-axis represents water level in a rice paddy, and $x$-axis shows growth stages. The growing season naturally separates into two phases: initial (left of dotted line) and main (right), which have distinct irrigation needs. Blue arrows indicate irrigation events, while yellow arrows denote sowing. The different practices create distinct temporal signatures, with the vertical dotted line marking the boundary between sowing and irrigation dimensions.
Figure 4: Dimensional divisions of dataset by practice type. A) Plots divided by sowing and irrigation dimensions, with number of plots in parentheses. B) Classification of irrigation practices (grouping Control and DSR as CF). C) Classification of sowing practices (grouping Control and AWD as PTR). Dotted red lines in B) and C) show classes are separated in a task-dependent manner.
Challenges
@@
Water Management Practices and Separability:
Water-saving practices in rice cultivation occur at different crop stages and create distinct water level patterns. DSR affects water management during early sowing and tillering phases, while AWD impacts water dynamics from tillering through harvest. These practices create unique temporal signatures in water levels and soil moisture that can be detected via SAR remote sensing (ijgi8070312; fikriyahDiscriminatingTransplantedDirect2019; villano2019separability; gumma2015mapping). As shown in Figure 4, PTR maintains consistently high water levels, while DSR shows a sharp initial decline after pre-sowing irrigation. Similarly, CF maintains flooded fields throughout the season, whereas AWD creates cyclical wet-dry patterns.
Dimensional Classification Approach:
@@
Planting Date Heterogeneity:
Satellite-based monitoring faces significant variation in planting schedules. Previous studies relied on synchronizing satellite data with known planting dates (fikriyahDiscriminatingTransplantedDirect2019; villano2019separability) – a critical challenge for monitoring millions of plots. In our study area, planting dates span up to 110 days, with DSR plots typically sown first, followed by control plots and AWD plots (Fig. S3).
Plot-level Predictions:
To provide fine-grained insights, we must be able to both identify individual plot boundaries as well as make predictions about these monolithic entities. We employ Fields of The World (FTW) algorithm (kerner2024fields) to automate the extraction of plot boundaries, enabling us to make predictions at an individual plot level.
@@ 3 Materials and Methods
Let X = $x_{t}]_{t=1}^{T}$ be the seasonal S1 feature sequence for a plot. We want to learn two predictive functions: sowing $f_{s}:X\rightarrow\hat{y}_{s}\in\{\text{DSR,PTR}\}$ and irrigation $f_{i}:X\rightarrow\hat{y}_{i}\in\{\text{AWD,CF}\}$ as parallel binary tasks, additionally fused into a joint prediction, $\hat{y}_{joint}=g(\hat{y}_{s},\hat{y}_{i})$. We aim to learn these predictors under limited ground truth and heterogeneous planting dates.
Ground Truth Data Collection
We partnered with the PRANA project to obtain reliable training data. Field teams collected detailed plot-level data from intervention sites across 18 districts in Punjab during the 2024 kharif rice season. The resulting dataset includes 452 AWD plots, 420 DSR plots and 411 PTR plots employing CF, with mean areas of 4.2, 2.7 and 2.1 acres, respectively (see Figure 2 for samples). Sowing and transplantation dates, irrigation schedules, and precise field boundaries were documented through on-site visits. Field polygons were digitized from 2024 Google Earth Pro imagery (30–50 cm) enforcing a minimum area of 2000 m2, and validated in-person for a subset of plots. A negative 10m buffer was applied to mitigate co-registration and boundary errors between the Google Earth basemap and co-registered S1 tiles.
Satellite Data and Processing
We leveraged S1 SAR data to monitor water management practices across the growing season (April - December 2024). We acquired 108 images from ascending orbit covering Punjab, focusing on May 1 - December 15222Through experimentation, the ascending orbit dataset was shown to have higher performance and was selected for training and inference. Previous studies have shown SAR’s effectiveness in capturing diverse rice agriculture metrics, including cultural type classification, phenological parameter detection, growing season estimation (clauss2018estimating; yangSystematicMethodSpatiotemporal2021; sahDiscriminationMonitoringRice2023), and differentiation between multiple cropping cycles and field-level tillage practices (singhaHighResolutionPaddy2019; liu2022using). SAR data was also selected for its cloud-penetrating capabilities during the monsoon season.
Features 3-Class 2-Class
Group HC DTW FT AE P #Feat. One-shot Fused Sowing Irrigation
Acc F1 Acc F1 Acc F1 Acc F1
Random $\times$ $\times$ $\times$ $\times$ $\times$ 0.333 0.332 0.333 0.334 0.559 0.559 0.542 0.540
Baselines $\checkmark$ 61 0.574 0.567 0.527* 0.534* 0.775* 0.769* 0.698* 0.687*
$\checkmark$ 3 0.450 0.450 0.434 0.431 0.636 0.639 0.612 0.612
$\checkmark$ 96 0.496 0.496 0.496 0.502 0.744 0.742 0.721* 0.707*
AE $\checkmark$ 64 0.450 0.450 0.450 0.453 0.721 0.715 0.635 0.625
$\checkmark$ $\checkmark$ 125 0.620 0.616 0.574 0.578 0.798 0.796 0.713* 0.697*
$\checkmark$ $\checkmark$ 160 0.589 0.588 0.597 0.601 0.783 0.774 0.744 0.742
Presto $\checkmark$ 6272 0.535* 0.526* 0.581 0.583 0.760 0.755 0.721 0.724
$\checkmark$ $\checkmark$ 6333 0.550 0.541 0.567 0.567 0.744 0.745 0.736 0.739
$\checkmark$ $\checkmark$ 6368 0.543* 0.534 0.589 0.589 0.752 0.750 0.721 0.725
AE + P $\checkmark$ $\checkmark$ 6336 0.581* 0.575* 0.581 0.583 0.752 0.750 0.736 0.736
$\checkmark$ $\checkmark$ $\checkmark$ 6432 0.581* 0.576* 0.605 0.603 0.752 0.750 0.729 0.734
Table 1: Classification performance by dimension of separation. Dimensional separation consistently outperforms 3-class classification for all tasks. Features: HC = hand crafted, DTW = Dynamic Time Weighting + kNN, FT = Fourier transform, P = Presto, AE = AlphaEarth. Results are from LightGBM unless *=RF. Bold indicates best results per task; underlined, second-best. All tasks use the best date interval from ablations, while AE features are fixed for the calendar year 2024. ‘One-shot’ refers to a 3-class classification which will mechanically have lower scores, and hence we compare to ‘fused’ which is the combination of making a DSR prediction first, followed by AWD for a given plot. Acc = accuracy, F1 = F1-weighted. #Feat. indicates the number of features used for that task.
@@
All imagery was processed using the SNAP toolbox (filipponi2019sentinel) to extract time series at 10-meter resolution. We extracted VV (Vertical transmit, Vertical receive) and VH (Vertical transmit, Horizontal receive) polarization bands, applying standard preprocessing: orbit file application, noise removal, radiometric calibration, multi-temporal speckle filtering (yommy2015sar), and terrain correction to obtain terrain-geocoded backscatter coefficient, $\gamma_{0}$. For each plot, we calculated mean backscatter values per timestep, generating plot-level timeseries. To handle temporal gaps, we applied temporal smoothing with a 1-D Gaussian filter followed by third-degree polynomial spline fitting, enabling regular sampling despite variable acquisition dates, crucial for dealing with missing timeseries data. We additionally create a VV/VH band ratio time series, which has proven effective for rice mapping in previous studies (villano2019separability; bazziMappingPaddyRice2019; singha2019high).
Feature Engineering
We derive complementary feature families that summarize both explicit temporal structure and learned representations of radar backscatter dynamics throughout the rice-growing season. All features are computed at the plot level from smoothed S1 VV, VH, and VV/VH ratio time series.
@@
Analytical Features: For each polarization (VV, VH, VV/VH), we extract temporal characteristics that include the position and amplitude of the first $k$ troughs, crests, and inflection points; the total number of troughs and crests; summary statistics (mean, minimum, maximum amplitude); parameters from Gaussian kernel fitting to the VV/VH ratio (amplitude, peak day, standard deviation and $R^{2}$(bazziMappingPaddyRice2019)); and radar vegetation index (RVI) values (nasirzadehdizaji2019sensitivity) for a total of 61 features, which we refer to as hand-crafted (HC) features. To complement these direct time-domain descriptors, we compute Fourier-domain features using fast Fourier transform (FT) that capture periodic irrigation-drainage patterns and overall signal smoothness.
Learnt Features: We include pretrained representations from two EO foundation models. The Pretrained Remote Sensing Transformer (Presto) (tseng2023lightweight) encodes up to 24 Sentinel-1 timesteps into 128-dimensional temporal embeddings that capture non-linear phenological trajectories. The Google AlphaEarth (AE) embeddings product (brown2025alphaearth) provides lightweight 64-dimensional embeddings summarizing annual multisensor (S1, S2, Landsat) variability, capturing landscape level dynamics at 10-meter resolutions. Both these embeddings are inexpensive to acquire, and therefore very attractive to practitioners.
Classification Approach
We test two ensemble learning methods—Random Forest (RF) and Gradient Boosted Trees (GB)–for their proven effectiveness in agricultural classification tasks (liuUsingSentinel1Sentinel22022; fangComprehensiveReviewRice2024a). These models effectively handle the heterogeneous feature sets while remaining computationally efficient. We also include a Dynamic Time Warping (DTW) baseline to benchmark performance without explicit feature extraction. DTW directly compares raw S1 time series using dynamic alignment and classifies them via a 1-nearest-neighbor scheme implemented using the tslearn package (zhang2017dynamic).
@@
Following our dimensional framework, we structure three classification tasks: a) AWD vs. DSR vs. Control (null hypothesis: all practices on same dimension), b) DSR vs. PTR (sowing dimension), and c) AWD vs. CF (irrigation dimension). For each task, we split data into non-overlapping train/test sets (90:10 ratio) using stratified sampling. We optimize hyperparameters using Optuna (akiba2019optuna) and evaluate performance using overall accuracy and F1 scores (globally weighted averages). We test a set of combinations of features and classification methods, summarized in Table 1. This approach tests our hypothesis that separating water management practices along their natural dimensions (sowing and irrigation) produces more accurate results than attempting simultaneous multi-class classification (i.e. the null hypothesis, 3-class one-shot). For comparison, we also evaluate a fused three-class setup where the sowing and irrigation outputs are combined into a single categorical label via rule-based fusion ($\text{fused}:=\text{DSR}\rightarrow\text{DSR}$, else irrigation decides AWD vs CF) - illustrating compounded error from sequential binary decisions and serves as a fairer baseline.All results use stratified random splits within the same region-year (Punjab-2024). Please see Figure A3 for the complete training workflow.
Temporal Ranges Sowing Irrigation F1 (weighted)
Row Dates M J J A S O N D Lag (%) DSR PTR CF AWD ALL DSR AWD
1 May 1 - Aug 15* $\sim$50% $\checkmark$ $\checkmark$ >> $\times$ 0.519 0.740 0.718
2 Jun 1 - Aug 30 $\sim$50% >> $\checkmark$ $\checkmark$ << 0.539 0.720 0.723
3 Jun 1 - Sep 5* $\sim$50% >> $\checkmark$ $\checkmark$ << 0.539 0.755 0.723
4 Jun1 - Oct 15 $\sim$60% >> $\checkmark$ $\checkmark$ $\checkmark$ 0.543 0.751 0.731
5 July 1 - Oct 15 $\sim$50% $\times$ $\checkmark$ $\checkmark$ $\checkmark$ 0.531 0.728 0.719
6 Aug 1 - Oct 15 $\sim$33% $\times$ $\times$ $\checkmark$ $\checkmark$ 0.489 0.696 0.709
7 Aug 1 - Nov 15 $\sim$50% $\times$ $\times$ $\checkmark$ $\checkmark$ 0.434 0.728 0.707
8 Sep 1 - Dec 15 $\sim$50% $\times$ $\times$ $\checkmark$ $\checkmark$ 0.499 0.677 0.662
9 Oct 1 - Dec 15 $\sim$33% $\times$ $\times$ << $\checkmark$ 0.519 0.658 0.643
10 Aug 1 - Dec 15 $\sim$60% $\times$ $\times$ $\checkmark$ $\checkmark$ 0.470 0.701 0.683
11 Jun 1 - Dec 15 $\sim$33% >> $\checkmark$ $\checkmark$ $\checkmark$ 0.541 0.713 0.737
12 May 1 - Dec 15** $\sim$90% $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 0.537 0.713 0.739
Table 2: Effect of temporal range selection on accuracy. Features are generated for different temporal ranges to analyze the effect on outcomes by manipulating the availability of lag and practice features for Presto+ features. AE features are not used, since its temporal range is fixed to a calendar year. Cell colors for months represent how much of the month is covered for that row (blank=none, green=full, orange=half). Values in bold indicate the top-2 performers for each classification task. Nominal sampling frequency of 7 days; *= a sampling frequency of 4 days, **=10 days. $\checkmark$ and $\times$ indicate the availability of a feature type, while >> & << imply that most & some features available, respectively. Sowing and irrigation feature availability is decided by virtue of which part of the growing cycle the temporal range reveals.